The 2014 balance sheet of Maria’s Tennis Shop, Inc., showed long-term debt of $5.3 million, and the 2015 balance sheet showed long-term debt of $6 million. The 2015 income statement showed an interest expense of $91026. What was the firm’s cash flow to creditors during 2015
Answer #1

Net new long term debt=Long term debt for 2015-Long term debt for 2014

=(6,000,000-5,300,000)=$700,000

Hence cash flow to creditors=Interest expense-Net new long term debt

=(91026-700,000)

=($608,974)(Negative).